OpenStudy (anonymous):

A satelite in circular orbit 1250 kilometers above Earth makes one complete revolution every 110 minutes. What is its linear speed? Assume that Earth is a sphere of radius 6400 kilometers

OpenStudy (anonymous):

step 1. the radius of the orbit for satellite is the sum of the radius of the earth and distance above earth so orbit radius is: R = 1250 + 6400 = 7650 km Time taken for one revolution T = 110 mins = 110/60 = 11/6 hours (1 hr = 60 mins) step 2: total linear distance covered in one revolution of a radius R orbit is S = 2*pi*R (its perimeter of the the circular orbit pi = 22/7 approx) S = 2*(22/7)*7650 km step 3: linear speed = Linear distance / Total time taken = 2*(22/7)*7650 / (11/6) km/hr linear speed = 26228.57 km/hr

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

Thank you! Having to take the 2nd semester of precal in order to graduate early in december without having been taught the first yet! I appreciate it.